Qd = 25000 – 2P

Qs =10000 + 1P

Calculate price elasticity of demand using point elasticity method when the construction industry is in equilibrium and interpret the result? 


1
Expert's answer
2021-05-09T14:29:08-0400

"Qd = Qs"

"25000-2p= 10000+1p"

"25000-10000=P+2P"

"15000=3P"  

"P= 5000"


"Qd = 25000-2p"  

"= 25000-2(5000)"

"=25000-10000"

"Q=15000"


Point elasticity of demand "= (\frac{∆Qd}{∆P})\times(\frac{P}{Q})"

"= -2\times(\frac{5000}{15000})"

"=\frac{-2}{3}"

"=0.67<1"

The elasticity of demand at equilibrium is inelastic. That is , when price increases, the quantity is not much affected at equilibrium point.